OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (4, 10)

OpenStudy (anonymous):

To write your equation in this way, y=mx+c, you don't know what are 'm' and'c' but you know what is 'x' and'y' To find 'm', you know this equation is parallel to y = x + 8. 'm' is gradient. 2 parallel lines have same gradient, so m=1 because the 'm' of y = x + 8 is 1. To find 'c' is very simple. 'c' is y intercept. Now you know what are 'x', 'y' and 'm'. Just substitute these value in and you can find 'c'. You can try it yourself and post your answer to check if it is right or not. I hope this helps.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

no problem